# Node Duplication Improves Cold-start Link Prediction

### v1

@@ 4.3 Performance Compare to Cold-start Methods
4.4 Performance Compared to Augmentation Methods
Effectiveness Comparison. Since NodeDup and NodeDup(L) use graph data augmentation techniques, we compare them to other data augmentation baselines. The performance and time consumption results are presented in Figure 5 for three datasets (Citeseer, Physics, and IGB-100K), while the results for the remaining datasets are provided in Appendix D due to the page limit. From Figure 5, we consistently observe that NodeDup outperforms all the graph augmentation baselines for Isolated and Low-degree nodes across all three datasets. Similarly, NodeDup(L) outperforms graph data augmentation baselines on 17/18 cases for Isolated and Low-degree nodes. Not only did our methods perform better for Isolated and Low-degree nodes, NodeDup and NodeDup(L) also perform on par or above baselines for Warm nodes.
Efficiency Comparison. Augmentation methods often come with the trade-off of adding additional run time before or during model training. For example, LAGNN (Liu et al., 2022b) requires extra preprocessing time to train the generative model prior to GNN training. It also takes additional time to generate extra features for each node during training. Although Dropedge (Rong et al., 2019) and TuneUP (Hu et al., 2022) are free of preprocessing, they require additional time to drop edges in each training epoch compared to base GNN training. Furthermore, the two-stage training employed by TuneUP doubles the training time compared to one-stage training methods. For NodeDup methods, duplicating nodes and adding edges is remarkably swift and consumes significantly less preprocessing time than other augmentation methods. As an example, NodeDup(L) and NodeDup are 977.0$\times$ and 488.5$\times$ faster than LAGNN in preprocessing Citeseer, respectively. We also observe that NodeDup(L) has the least training time among all augmentation methods and datasets, while NodeDup also requires less training time in 8/9 cases. Additionally, NodeDup(L) achieves significant efficiency benefits compared to NodeDup in Figure 5, especially when the number of nodes in the graph increases substantially. Taking the IGB-100K dataset as an example, NodeDup(L) is 1.3$\times$ faster than NodeDup for the entire training process.
4.5 Performance under the Inductive Setting
Under the inductive setting (Guo et al., 2022b; Shiao et al., 2022), which closely resembles real-world LP scenarios, the presence of new nodes after the training stage adds an additional challenge compared to the transductive setting. We evaluate and present the effectiveness of our methods under this setting in Table 2 for Citeseer, Physics, and IGB-100K datasets. Additional results for other datasets can be found in Appendix D. In Table 2, we observe that our methods consistently outperform base GSage across all of the datasets. We also observe significant improvements in the performance of our methods on Isolated nodes, where NodeDup and NodeDup(L) achieve 5.50% and 3.57% improvements averaged across the three datasets, respectively. Additionally, NodeDup achieves 5.09% improvements on Low-degree nodes. NodeDup leads to more pronounced improvements on Low-degree/Isolated nodes, making it particularly beneficial for the inductive setting.
4.6 Performance with Different Encoders/Decoders
As a simple plug-and-play augmentation method, NodeDup can work with different GNN encoders and LP decoders. In Tables 3 and 4, we present results with GAT (Veličković et al., 2017) and JKNet (Xu et al., 2018) as encoders, along with a MLP decoder. Due to the space limit, we only report the results of three datasets here and leave the remaining in Appendix D. When applying NodeDup to base LP training, with GAT or JKNet as the encoder and inner product as the decoder, we observe significant performance improvements across the board. Regardless of the encoder choice, NodeDup consistently outperforms the base models, particularly for Isolated and Low-degree nodes. Meanwhile, we note that the improvement achieved with our methods is more pronounced when using JKNet as the encoder compared to GAT. This can be attributed to the more informative multi-view provided by JKNet during the aggregation step, as discussed in Section 3.2.
Table 4: LP performance with MLP decoder (GSage as the encoder). Our methods perform better than the base model.
@@ 4.6 Performance with Different Encoders/Decoders
In Table 4, we present the results of our methods applied to the base LP training, where GSage serves as the encoder and MLP as the decoder. Regardless of the decoder, we observe better performance with our methods. These improvements are significantly higher compared to the improvements observed with the inner product decoder. The primary reason for this discrepancy is the inclusion of additional supervised training signals for isolated nodes in our methods, as discussed in Section 3.2. These signals play a crucial role in training the MLP decoder, making it more responsive to the specific challenges presented by isolated nodes.
@@ 5 Conclusion
Impact Statement. In this work, our simple but effective method enhances the link prediction performance on cold-start nodes, which mitigates the degree bias and advances the fairness of graph machine learning. It can be widely used and beneficial for various real-world applications, such as recommendation systems, social network analysis, and bioinformatics. We do not foresee any negative societal impact or ethical concerns posed by our method. Nonetheless, we note that both positive and negative societal impacts can be made by applications of graph machine learning techniques, which may benefit from the improvements induced by our work. Care must be taken, in general, to ensure positive societal and ethical consequences of machine learning.
@@ Appendix A Related Work
LP with GNNs. Over the past few years, GNN architectures (Kipf & Welling, 2016a; Gilmer et al., 2017; Hamilton et al., 2017; Veličković et al., 2017; Xu et al., 2018) have gained significant attention and demonstrated promising outcomes in LP tasks. There are two primary approaches to applying GNNs in LP. The first approach involves a node-wise encoder-decoder framework, which we discussed in Section 2. The second approach reformulates LP tasks as enclosing subgraph classification tasks (Zhang & Chen, 2018; Cai & Ji, 2020; Cai et al., 2021; Dong et al., 2022). Instead of directly predicting links, these methods perform graph classification tasks on the enclosing subgraphs sampled around the target link. These methods can achieve even better results compared to node-wise encoder-decoder frameworks by assigning node labels to indicate different roles within the subgraphs. However, constructing subgraphs poses challenges in terms of efficiency and scalability, requiring substantial computational resources. Our work focuses on the encoder-decoder framework for LP, circumventing the issues associated with subgraph construction.
Methods for Cold-start Nodes. Recently, several GNN-based methods (Wu et al., 2019; Liu et al., 2020; Tang et al., 2020b; Liu et al., 2021; Zheng et al., 2021) have explored degree-specific transformations to address robustness and cold-start node issues. Tang et al.(Tang et al., 2020b) introduced a degree-related graph convolutional network to mitigate degree-related bias in node classification tasks. Liu et al.(Liu et al., 2021) proposed a transferable neighborhood translation model to address missing neighbors for cold-start nodes. Zheng et al.(Zheng et al., 2021) tackled the cold-start nodes problem by recovering missing latent neighbor information. These methods require cold-start-node-specific architectural components, unlike our approach, which does not necessitate any architectural modifications. Additionally, other studies have focused on long-tail scenarios in various domains, such as cold-start recommendation(Chen et al., 2020; Lu et al., 2020; Hao et al., 2021). Imbalance tasks present another common long-tail problem, where there are long-tail instances within small classes (Lin et al., 2017; Ren et al., 2020; Tan et al., 2020; Kang et al., 2019; Tang et al., 2020a). Approaches like (Lin et al., 2017; Ren et al., 2020; Tan et al., 2020) address this issue by adapting the loss for different samples. However, due to the different problem settings, it is challenging to directly apply these methods to our tasks. We only incorporate the balanced cross entropy introduced by Lin et al. (Lin et al., 2017) as one of our baselines.
@@ Appendix B Additional Datasets Details
Why there are no OGB (Hu et al., 2020) datasets applied? OGB benchmarks that come with node features, such as OGB-collab and OGB-citation2, lack a substantial number of isolated or low-degree nodes, which makes it challenging to yield convincing results for experiments focusing on the cold-start problem. This is primarily due to the split setting adopted by OGB, where the evaluation is centered around a set of the most recent papers with high degrees. Besides, considering these datasets have their fixed splitting settings based on time, it will lead to inconsistent problems to compared with the leaderboard results if we use our own splitting method to ensure we have a reasonable number of isolated/low-degree nodes. Given these constraints, we opted for another extensive benchmark dataset, IGB-100K (Khatua et al., 2023), to test and showcase the effectiveness of our methods on large-scale graphs. We further conducted the experiments on IGB-1M, which are shown in Section D.8.
@@ B.1 Transductive Setting
For the transductive setting, we randomly split the edges into training, validation, and testing sets based on the splitting ratio specified in Section 4.1. The nodes in training/validation/testing are all visible during the training process. However, the positive edges in validation/testing sets are masked out for training. After the split, we calculate the degrees of each node using the validation graph. The dataset statistics are shown in Table 5.
@@ B.2 Inductive Setting
The inductive setting is considered a more realistic setting compared to the transductive setting, where new nodes appear after the training process. Following the inductive setting introduced in Guo et al. (2022b) and Shiao et al. (2022), we perform node splitting to randomly sample 10% nodes from the original graph as the new nodes appear after the training process. The remaining nodes are considered observed nodes during the training. Next, we group the edges into three sets: observed-observed, observed-new, and new-new node pairs. We select 10% of observed-observed, 10% of observed-new, and 10% of new-new node pairs as the testing edges. We consider the remaining observed-new and new-new node pairs, along with an additional 10% of observed-observed node pairs, as the newly visible edges for the testing inference. The datasets statistics are shown in Table 5.
Appendix C NodeDup Algorithm
In this section, we provide a detailed description of our algorithm, which is outlined in Algorithm 1. Compared to the default training of GNNs for LP tasks, NodeDup incorporates additional augmentation steps, denoted as L1-L5 in Algorithm 1.
Algorithm 1 NodeDup.
[…]
2:  Identify cold node set $\mathcal{V}_{cold}$ based on the node degree.
3:  Duplicate all cold nodes to generate the augmented node set $\mathcal{V}^{\prime}=\mathcal{V}\cup\mathcal{V}_{cold}$, whose node feature matrix is then $\mathbf{X}^{\prime}\in\mathbb{R}^{(N+|\mathcal{V}_{cold}|)\times F}$.
4:  Add an edge between each cold node $v\in\mathcal{V}_{cold}$ and its duplication $v^{\prime}$, then get the augmented edge set =$\mathcal{E}^{\prime}=\mathcal{E}\cup\{e_{vv^{\prime}}:\forall v\in\mathcal{V}_{cold}\}$.
5:  Add the augmented edges into the training set and get $\mathcal{Y}^{\prime}=\mathcal{Y}\cup\{y_{vv^{\prime}}=1:\forall v\in\mathcal{V}_{cold}\}$.
6:  # End-to-end supervised training based on the augmented graph $G^{\prime}=\{\mathcal{V}^{\prime},\mathcal{E}^{\prime},\textbf{X}^{\prime}\}$.
7:  for $l=1$ to $L$ do
8:    for $v$ in $\mathcal{V^{\prime}}$ do
10:    end for
11:  end for
12:  for $(i,j)$ in $\mathcal{Y^{\prime}}$ do
14:  end for
Appendix D Further Experimental Results
D.1 Selection of the threshold $\delta$.
Our decision to set the threshold $\delta$ at 2 is grounded in data-driven analysis, as illustrated in Figure 1 and Figure 6. These figures reveal that nodes with degrees not exceeding 2 consistently perform below the average Hits@10 across all datasets, and higher than 2 will outperform the average results. Besides, our choice aligns with methodologies in previous studies (Liu et al., 2020, 2021), where cold nodes are identified using a fixed threshold across all the datasets. In addition, we conduct experiments with different thresholds $\delta$ on Cora and Citeseer datasets. The results are shown in Table 6. Our findings were consistent across different thresholds, with similar observations at $\delta$ = 1, $\delta$ = 2, and $\delta$ = 3. This indicates that our method’s effectiveness is not significantly impacted by changes in this threshold.

### v2

@@ 4.4 Performance Compared to Augmentation Methods
4.5 Performance under the Inductive Setting
Under the inductive setting (Guo et al., 2022; Shiao et al., 2022), which closely resembles real-world LP scenarios, the presence of new nodes after the training stage adds an additional challenge compared to the transductive setting. We evaluate and present the effectiveness of our methods under this setting in Table˜2 for Citeseer, Physics, and IGB-100K datasets. Additional results for other datasets can be found in Section˜C.10. In Table˜2, we observe that our methods consistently outperform base GSage across all of the datasets. We also observe significant performance improvements of our methods on Isolated nodes, where NodeDup and NodeDup(L) achieve 5.50% and 3.57% improvements averaged across the three datasets, respectively. Additionally, NodeDup achieves 5.09% improvements on Low-degree nodes. NodeDup leads to more pronounced improvements on Low-degree/Isolated nodes, making it particularly beneficial for the inductive setting.
4.6 Performance with Different Encoders/Decoders
As a simple plug-and-play augmentation method, NodeDup can work with different GNN encoders and LP decoders. In Tables˜3 and 4, we present results with GAT (Veličković et al., 2017) and JKNet (Xu et al., 2018) as encoders, along with a MLP decoder. Due to the space limit, we only report the results of three datasets here and leave the remaining in Section˜C.11. When applying NodeDup to base LP training, with GAT or JKNet as the encoder and inner product as the decoder, we observe significant performance improvements across the board. Regardless of the encoder choice, NodeDup consistently outperforms the base models, particularly for Isolated and Low-degree nodes. From Section˜C.11, we also observe the performance improvements of NodeDup with GCN (Kipf & Welling, 2016a), GraphTransformer (Dwivedi & Bresson, 2020) as encoders.
@@ 4.6 Performance with Different Encoders/Decoders
Table 4: LP performance with MLP decoder (GSage as the encoder). Our methods outperform the base model.
@@
In Table˜4, we present the results of our methods applied to the base LP training, where GSage serves as the encoder and MLP as the decoder. Regardless of the decoder, we observe better performance with our methods. These improvements are significantly higher compared to the improvements observed with the inner product decoder. The primary reason for this discrepancy is the inclusion of additional supervised training signals for isolated nodes in our methods, as discussed in Section˜3.2. These signals play a crucial role in training the MLP decoder, making it more responsive to the specific challenges presented by isolated nodes.
Furthermore, we apply NodeDup to SEAL (Zhang & Chen, 2018), a subgraph-based LP model, and observe notable performance gains, as shown in Section˜C.11, demonstrating the broad applicability of our method across different GNN models and training paradigms.
4.7 Influence of the Duplication Frequency and Duplicated Node Types in NodeDup
In our experiments, we duplicate the cold nodes once and add one edge for each cold node in NodeDup. Figure˜9 presents the results of our ablation study on Citeseer dataset, which examines how varying the duplication frequency and the types of duplicated nodes affect the performance of NodeDup across the Isolated, Low-degree, and Overall settings. The numbers in each block represent the performance differences relative to the baseline of duplicating cold nodes once. From the results, we observe that increasing the duplication frequency does not consistently lead to performance improvements across all settings. Notably, duplicating all nodes multiple times significantly enhances the performance of Isolated nodes. However, this approach also introduces a large number of isolated nodes into the graph, which negatively impacts the overall performance. Consequently, duplicating only cold nodes once emerges as the most effective strategy, as it consistently maintains strong performance across all settings. Further analysis of the impact of different duplicated node types is provided in Section˜C.2.
(a) Isolated
(b) Low-degree
(c) Overall
Figure 9: Ablation study on duplication frequency and node types of NodeDup. (a), (b), (c) show the performance for Isolated nodes, Low-degree nodes, and Overall settings, respectively. The values in each block represent the performance differences compared to the baseline setting of duplicating cold nodes once.
4.8 Analyzing Performance Gains of NodeDup on Warm Nodes
Table 5: Distribution and AUC performance of testing Warm-Warm and Warm-Cold links. NodeDup improves Warm-Cold link performance while maintaining Warm-Warm link performance.
Warm-Warm Warm-Cold
Number GSage NodeDup(L) NodeDup Number GSage NodeDup(L) NodeDup
Cora 157738 94.92$\pm$0.31 95.17$\pm$0.19 95.18$\pm$0.18 16759 77.06$\pm$1.40 81.41$\pm$1.18 80.51$\pm$1.72
Citeseer 63266 97.21$\pm$0.09 97.06$\pm$0.21 97.02$\pm$0.12 24020 85.40$\pm$0.78 87.96$\pm$0.79 88.40$\pm$0.92
CS 4209161 98.31$\pm$0.03 98.30$\pm$0.02 98.42$\pm$0.02 91458 87.92$\pm$0.19 91.47$\pm$0.35 90.44$\pm$0.84
Physics 11462743 99.01$\pm$0.01 99.01$\pm$0.02 99.02$\pm$0.00 103174 86.21$\pm$0.33 89.94$\pm$0.31 90.23$\pm$0.51
Photos 2984253 97.85$\pm$0.06 98.03$\pm$0.04 97.87$\pm$0.02 104737 59.80$\pm$1.33 68.11$\pm$0.43 64.32$\pm$0.73
Computers 5417165 97.58$\pm$0.07 97.60$\pm$0.08 97.54$\pm$0.09 217090 46.49$\pm$0.75 57.32$\pm$0.99 57.63$\pm$0.49
IGB-100K 6899924 98.70$\pm$0.00 98.71$\pm$0.02 98.64$\pm$0.01 1372994 97.14$\pm$0.10 98.63$\pm$0.42 98.23$\pm$0.06
To better understand the performance improvements of NodeDup on Warm nodes, we first analyzed the distribution of Warm-Warm and Warm-Cold links in the testing set. Our analysis reveals that, across all datasets, the number of Warm-Warm links consistently exceeds that of Warm-Cold links, as shown in Table˜5. This observation indicates that Warm-Cold links do not overwhelmingly influence the overall performance of Warm nodes. To further investigate, we conducted experiments comparing the performance of our method on these two types of links. The results demonstrate that our approach consistently enhances performance on Warm-Cold links while maintaining strong performance on Warm-Warm links. These findings confirm that our method does not negatively impact the learning ability of Warm nodes. Instead, the observed performance gains primarily stem from improved learning on Cold nodes, as previously discussed in Section˜4.2.
@@ 5 Conclusion
Acknowledgements
We are grateful for all the insightful comments and constructive suggestions provided by the reviewers. This research was partially conducted during Zhichun’s internship at Snap Inc. and was supported by the National Science Foundation (NSF) through the Center for Computer Assisted Synthesis (C-CAS), under grant number CHE-2202693.
@@ Appendix A Related Work
LP with GNNs. Over the past few years, GNN architectures (Kipf & Welling, 2016a; Gilmer et al., 2017; Hamilton et al., 2017; Veličković et al., 2017; Xu et al., 2018) have gained significant attention and demonstrated promising outcomes in LP tasks. There are two primary approaches to applying GNNs in LP. The first approach involves a node-wise encoder-decoder framework, which we discussed in Section˜2. The second approach reformulates LP tasks as enclosing subgraph classification tasks (Zhang & Chen, 2018; Cai & Ji, 2020; Cai et al., 2021; Dong et al., 2022). Instead of directly predicting links, these methods perform graph classification tasks on the enclosing subgraphs sampled around the target link. These methods can achieve even better results compared to node-wise encoder-decoder frameworks by assigning node labels to indicate different roles within the subgraphs. However, constructing subgraphs poses challenges in terms of efficiency and scalability, requiring substantial computational resources. Our work focuses on the encoder-decoder framework for LP, circumventing the issues associated with subgraph construction.
Methods for Cold-start Nodes. Recently, several GNN-based methods (Wu et al., 2019; Liu et al., 2020; Tang et al., 2020b; Liu et al., 2021; Zheng et al., 2021) have explored degree-specific transformations to address robustness and cold-start node issues. Tang et al.(Tang et al., 2020b) introduced a degree-related graph convolutional network to mitigate degree-related bias in node classification tasks. Liu et al.(Liu et al., 2021) proposed a transferable neighborhood translation model to address missing neighbors for cold-start nodes. Zheng et al.(Zheng et al., 2021) tackled the cold-start nodes problem by recovering missing latent neighbor information. These methods require cold-start-node-specific architectural components, unlike our approach, which does not necessitate any architectural modifications. Additionally, other studies have focused on long-tail scenarios in various domains, such as cold-start recommendation(Chen et al., 2020; Lu et al., 2020; Hao et al., 2021). Imbalance tasks present another common long-tail problem, where there are long-tail instances within small classes (Lin et al., 2017; Ren et al., 2020; Tan et al., 2020; Kang et al., 2019; Tang et al., 2020a). Approaches like (Lin et al., 2017; Ren et al., 2020; Tan et al., 2020) address this issue by adapting the loss for different samples. However, due to the different problem settings, it is challenging to directly apply these methods to our tasks. We only incorporate the balanced cross entropy introduced by Lin et al. (Lin et al., 2017) as one of our baselines. In addition to these one-shot training methods, many recommendation studies (Kirkpatrick et al., 2017; Xu et al., 2020; Zhou & Cao, 2021; Valkanas et al., 2024) have addressed the issue of incorporating new nodes through incremental learning frameworks, where new nodes are continually added into the training process and specific strategies are designed to alleviate forgetting previous knowledge (i.e., warm nodes or old classes).
@@ Appendix B Additional Datasets Details
Why there are no OGB (Hu et al., 2020) datasets applied? OGB benchmarks that come with node features, such as OGB-collab and OGB-citation2, lack a substantial number of isolated or low-degree nodes, which makes it challenging to yield convincing results for experiments focusing on the cold-start problem. This is primarily due to the split setting adopted by OGB, where the evaluation is centered around a set of the most recent papers with high degrees. Besides, considering these datasets have their fixed splitting settings based on time, it will lead to inconsistent problems to compared with the leaderboard results if we use our own splitting method to ensure we have a reasonable number of isolated/low-degree nodes. Given these constraints, we opted for another extensive benchmark dataset, IGB-100K (Khatua et al., 2023), to test and showcase the effectiveness of our methods on large-scale graphs. We further conducted the experiments on IGB1M, which are shown in Section˜C.3.
@@ B.1 Transductive Setting
For the transductive setting, we randomly split the edges into training, validation, and testing sets based on the splitting ratio specified in Section˜4.1. The nodes in training/validation/testing are all visible during the training process. However, the positive edges in validation/testing sets are masked out for training. After the split, we calculate the degrees of each node using the validation graph. The dataset statistics are shown in Table˜6.
@@ B.2 Inductive Setting
The inductive setting is considered a more realistic setting compared to the transductive setting, where new nodes appear after the training process. Following the inductive setting introduced in Guo et al. (2022) and Shiao et al. (2022), we perform node splitting to randomly sample 10% nodes from the original graph as the new nodes appear after the training process. The remaining nodes are considered observed nodes during the training. Next, we group the edges into three sets: observed-observed, observed-new, and new-new node pairs. We select 10% of observed-observed, 10% of observed-new, and 10% of new-new node pairs as the testing edges. We consider the remaining observed-new and new-new node pairs, along with an additional 10% of observed-observed node pairs, as the newly visible edges for the testing inference. The datasets statistics are shown in Table˜6.
Appendix C Further Experimental Results
C.1 Selection of the threshold $\delta$.
Our decision to set the threshold $\delta$ at 2 is grounded in data-driven analysis, as illustrated in Figure˜1 and Figure˜10. These figures reveal that nodes with degrees not exceeding 2 consistently perform below the average Hits@10 across all datasets, and higher than 2 will outperform the average results. Besides, our choice aligns with methodologies in previous studies (Liu et al., 2020; 2021), where cold nodes are identified using a fixed threshold across all the datasets. In addition, we conduct experiments with different thresholds $\delta$ on Cora and Citeseer datasets. The results are shown in Table˜7. Our findings were consistent across different thresholds, with similar observations at $\delta$ = 1, $\delta$ = 2, and $\delta$ = 3. This indicates that our method’s effectiveness is not significantly impacted by changes in this threshold.